ABC Company is planning to purchase an equipment. The purchase price of the equipment is $350,000. The company plans to make a down payment of 25% of the first cost, and for the remainder of the cost of the equipment, they plan to take a loan. The company will pay off this loan in 7 years at 10% in equal annual payments.

ABC believes that the equipment can be sold for $75,000 at the end of its 15-year service life. The new equipment will increase the company's annual income by $90,000. Maintenance and operating costs are expected to be $4,000 during the first year and to increase by $1,200 each year. ABC uses a before-tax MARR of 12% for its preliminary economic studies. Determine

i) Annual payment amount for the loan. (15 points)

ii) Present worth of costs (down & loan payments, maintenance & operation costs) at MARR. (24 points)

iii) Present worth of benefits (annual income, salvage value). (20 points)

iv) NPW of the investment to ABC. Is the investment desirable?
Answer #1

Initial Cost=$350000(1-0.75)=0.25*350000=$87500

Salvage Value=$75000

Annual Income=$90000

Annual M&O Cost=$4000 and increases by $1200 thereafter

MARR=12%

Lets first find Annual payment amount for the loan
=Principle*Rate*(1+R)^n/((1+R)^n-1)

Annual Payment amount for the loan=262500*0.1*(1.1)^7/((1.1)^7-1)=$53,918.94

Ans 2)

PW of Costs= Initial Cost+PW of Annual Loan payments+ PW of M&O costs

=87500+53918.94(P/A,12%,7)+4000(P/A,10%,7)+1200(P/G,10%,7)
=87500+53918.94(4.564)+4000(4.564)+1200(11.644)
PW of Costs=$365814.84

Ans 3)

PW of Benefits=PW of Annual Income+ PW of Salvage Value

=90000(P/A,12%,7)+75000(P/F,12%,7)

=90000(4.564)+75000(0.4523)

PW of Benefits=$444682.5

Ans 4)

NPW= PW of Benefits-PW of Costs

=444682.5-365814.84=$78867.66>0

As Net present worth is positive this investment is desirable.